OpenStudy (anonymous):

Find the value of y for a given value of x, if y varies directly with x. If y = 6 when x = 2, what is y when x = 5?

OpenStudy (anonymous):

okk so for this.... \[\frac{ x_{1} }{ y_{1} }=\frac{ x_{2} }{ y_{2} }\]

OpenStudy (anonymous):

2/6 = 5/y. "directly proportional" or "directly varies" means x/y is proportional to another x/y, so just make them equal. solve for y

OpenStudy (anonymous):

so then u get...... \[\frac{2}{6}=\frac{5}{y}\]

OpenStudy (anonymous):

now just solve for y

OpenStudy (anonymous):

how do i do that??

OpenStudy (anonymous):

would it be 5/15 ??

OpenStudy (anonymous):

yupp so 15 is your answer